40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Sat Nov 12 01:09:56 2016, gautam, Update, LSC, Recovering DRMI locking 
    Reply  Wed Nov 16 03:10:01 2016, gautam, Update, LSC, DRMI locked on 1f and 3f signals 
       Reply  Wed Nov 16 08:14:43 2016, Steve, Update, LSC, DRMI locked on 1f and 3f signals 5hrs.png
       Reply  Mon Nov 21 14:02:32 2016, gautam, Update, LSC, DRMI locked on 3f signals, arms held on ALS DRMIArms_Nov20.pdf
    Reply  Wed Nov 23 16:21:02 2016, gautam, Update, LSC, ITMY UL glitches are back ITMYproblemsBack.png
       Reply  Mon Nov 28 10:27:13 2016, gautam, Update, SUS, ITMY UL glitches are back ITMY_satboxSpectra.pdfITMY_UL_testerBox.png
       Reply  Tue Nov 29 11:07:37 2016, Steve, Update, LSC, ITMY UL glitches are back gliching400d.png
          Reply  Wed Nov 30 01:47:56 2016, gautam, Update, LSC, Suspension woes IMCwoes.png
             Reply  Wed Nov 30 17:08:56 2016, gautam, Update, LSC, Binary output breakout box removed 
                Reply  Thu Dec 1 02:19:13 2016, gautam, Update, LSC, Binary output breakout box restored MC1_glitch_fast.pngITMY_testerGlitch.pngITMYsatbox.pngULcomparison.pdf
                   Reply  Thu Dec 1 08:02:57 2016, Steve, Update, LSC, glitching ITMY_UL_LL ITMY_UL_LL.png
                      Reply  Fri Dec 2 11:56:42 2016, gautam, Update, LSC, MC1 LEMO jiggled IMG_3474.JPGMC1_LR_dead.png
                         Reply  Mon Dec 5 15:05:37 2016, gautam, Update, LSC, MC1 glitches are back Mc1glitches.png
                      Reply  Thu Dec 8 10:13:43 2016, Steve, Update, LSC, glitching ITMY_UL has a history glitching__ITMY-UL_2007.png
Message ID: 12611     Entry time: Sat Nov 12 01:09:56 2016     Reply to this: 12619   12638
Author: gautam 
Type: Update 
Category: LSC 
Subject: Recovering DRMI locking 

Now that we have all Satellite boxes working again, I've been working on trying to recover the DRMI 1f locking over the last couple of days, in preparation for getting back to DRFPMI locking. Given that the AS light levels have changed, I had to change the whitening gains on the AS55 and AS110 channels to take this into account. I found that I also had to tune a number of demod phases to get the lock going. I had some success with the locks tonight, but noticed that the lock would be lost when the MICH/SRCL boosts were triggered ON - when I turned off the triggering for these, the lock would hold for ~1min, but I couldn't get a loop shape measurement in tonight.


As an aside, we have noticed in the last couple of months glitchy behaviour in the ITMY UL shadow sensor PD output - qualitatively, these were similar to what was seen in the PRM sat. box, and since I was able to get that working again, I did a similar analysis on the ITMY sat. box today with the help of Ben's tester box. However, I found nothing obviously wrong, as I did for the PRM sat. box. Looking back at the trend, the glitchy behaviour seems to have stopped some days ago, the UL channel has been well behaved over the last week. Not sure what has changed, but we should keep an eye on this...

ELOG V3.1.3-